# Gravitomagnetism from Temporal Dimensional Reduction

### v4

@@ Abstract
We dimensionally reduce the Taub-NUT metric to three spatial dimensions by treating time as the extra curled dimension. Since closed timelike curves can exist in the Taub-NUT framework, such a dimensional reduction is justified. Thus, the results end up with $3$-dimensional corresponding Einstein field equations plus corresponding Maxwell type equations for a gravitomagnetic field, associated with the NUT charge, which also acts as a source for the Einstein field equations. In this approach, the Taub-NUT metric can be envisaged as a $(1+3)$-dimensional analogue of the $(1+4)$-dimensional metric of the Kaluza-Klein theory. Hence, in four dimensions, it unifies gravitation and gravitomagnetism, associated with the NUT charge, in the same footing that the Kaluza-Klein theory unifies gravitation and electromagnetism in five dimensions. In fact, in this way, a new gravitomagnetic field emerges from the temporal dimensional reduction. In another word, when the time coordinate is considered curled to a non-observable size, gravitation and gravitomagnetism, associated with the NUT charge, appear as two distinct fields. We also introduce a relation between the $4$-dimensional gravitational constant and the NUT charge.
@@ 1 Introduction
On the other hand, the Taub-NUT (Newman-Tamburino-Unti) solution [18, 19] to the Einstein equations of GR has been extensively studied in the context of gravitational physics and mathematical relativity [20][22]. One the interesting aspects of this solution is the existence of the NUT charge, which is the source of the gravitomagnetic field [^4] Such a gravitomagnetic field is different from the one resulting from the weak-field and slow-motion limit of GR [23], and was experimentally confirmed in $2011$ [24]. due to the gravitomagnetic monopoles [25][27], the gravitational analogue of the magnetic monopoles [28]. However, a problem with the Taub-NUT solution is the string singularity called the Misner string. One way to eliminate the Misner string is to make the time coordinate periodic [29], which leads to the existence of closed time-like curves (CTCs). CTCs are a hallmark of the Taub-NUT solution and have profound implications for causality in GR. A CTC is a path through spacetime that loops back on itself while remaining time-like throughout. Essentially, it allows for a type of time travel in which the traveler can revisit their past, which violates causality. Causality is a foundational concept in physics, ensuring that in all frames of reference, causes precede effects. The existence of CTCs challenges this principle by allowing situations in which the distinction between cause and effect is broken down. Hence, such a paradoxical behavior implies that the usual interpretation of spacetime as a causal structure, in the presence of CTCs, needs to be modified or scrutinized [30].
However, CTCs are not unique to the Taub-NUT metric, but appear in several solutions of the Einstein field equations, particularly in spacetimes with unusual topology or specific symmetry properties. Notable examples include: The Gödel solution as a rotational cosmological model that admits CTCs [31] and violates global causality everywhere. Also, the interior region (inside the event horizon) of a rotating black hole, as indicated by the Kerr and Kerr-Newman metrics [32], can host CTCs, particularly near the ring singularity. These CTCs are confined to the interior of black holes and do not affect causality in the external spacetime [32]. Traversable wormholes can create CTCs if their ends experience different time dilation effects [33]. Certain configurations of rotating cosmic strings can admit CTCs due to frame-dragging effects [34]. Cylindrical distributions of rotating dust, called the van Stockum dust, can produce CTCs, depending on the density and angular velocity of the dust [35].
The existence of CTCs in various solutions of the Einstein equations has sparked significant debate and research. The causality violations are often considered as artifacts of mathematical structure rather than realizable physical scenarios and numerous approaches have been adapted in the literature to modify spacetime metrics and eliminate CTCs. Some treatments suggest removing or modifying problematic regions of spacetime, such as cutting out those responsible for CTCs [20]. Also, Stephen Hawking proposed the ‘chronology protection conjecture’, hypothesizing that quantum effects prevent the formation of CTCs in physically realistic spacetimes [30]. Hawking famously said there seems to be a chronology protection agency that prevents the emergence of CTCs, thus making the universe safe for historians. Also, adjusting the boundary conditions at infinity is a way to avoid CTCs [29]. Nonetheless, such approaches to eliminate CTCs in the $4$-dimensional Taub-NUT metric come with trade-offs such as restricting spatial coordinates [20].
@@ 2 Brief Review of the Kaluza-Klein Theory
$S^{\rm(KK)}\equiv\hat{S}=\frac{1}{{\kappa^{2}}}\int d^{4}x\,\sqrt{-g}\,\phi\,\left(R-\frac{1}{4}\phi^{2}\kappa^{2}F_{\alpha\beta}F^{\alpha\beta}+2\,\frac{\partial^{\alpha}\phi\,\partial_{\alpha}\phi}{\phi^{2}}\right),$ (5)
@@ 3 The Taub-NUT Solution
The main significance of the Taub-NUT solutions is the existence of the Misner string at $\theta=0$ or $\theta=\pi$, which makes the $1$-form $d\varphi$ ill-defined at either of these points. To get rid of the singularity at $\theta=0$, one can shift the time coordinate such that $t^{\prime}=t+2n\varphi$, and for the point $\theta=\pi$, one can take $t^{\prime\prime}=t-2n\varphi$, where $n$ is an integer. Since $\varphi$ is periodic by $2\pi$, one has to identify $t^{\prime}$ and $t^{\prime\prime}$ by modulo $8\pi n$, which boils down to imposing the periodicity to the time coordinate as $t\rightarrow t+8\pi n$. Doing this makes the $1$-form $d\varphi$ well-defined but leads to CTCs in the Lorentzian metric. Importantly, the elimination of the Misner string also leads to a quantization condition, namely the quantization of the NUT charge [41], which also acts as a topological charge by determining the topological twisting of spacetime [42].
@@ 4 Time as an Extra Dimension in Taub-NUT Metric
To address the CTC issue, in this section we try to establish a connection between the Taub-NUT solution and the KK framework. The main purpose is to start from the Taub-NUT metric in a $4$-dimensional spacetime coordinate, when $l\neq 0$, and eventually arrive at a $3$-dimensional spatial metric plus a compactified or curled time dimension. Actually, we assume that there is a dynamical mechanism that drags the size of time dimension down and keeps it in a stable way at an unobservably tiny scale. To begin, we rewrite the $4$-dimensional Taub-NUT metric (9) in the matrix form [^9] We represent $3$-dimensional quantities with a tilde, $\widetilde{}$ , over them.
@@
where the lowercase Latin indices run from one to three, the solution to the $3$-dimensional vector potential is obtained as $\tilde{\bf A}=(0,0,\cos{\theta})$, the metric signature is (,+,+,+)(-,+,+,+) provided that $\tilde{g}_{33}-4l^{2}\Phi^{2}(r)\cos^{2}{\theta}>0$, and $g=-\Phi^{2}(r)\,\tilde{g}$, where $\tilde{g}$ is the determinant of the $3$-dimensional metric. Moreover, the solution to the $3$-dimensional metric is
@@
with the signature $(+,+,+)$ in the spherical coordinates.
@@
where $\oint dx^{0}$} is the result of integration along the $x^{0}$-direction. From the $3$-dimensional point of view, $\oint dx^{0}$} can be considered as the size of the CTC. The reduced action involves the gauge field $\tilde{A}_{a}$, which is a gravitomagnetic type gauge field associated with the NUT charge, and $\Phi(r)$, which plays the role of a scalar or a dilaton field within the context of the KK theory.
@@
Now, to see how the model deals with the CTC problem, we note that for $r_{-}<r<r_{+}$ including $r=l$, the scalar field $\Phi(r)$ (10) becomes imaginary and hence the signature of metric (9) changes to $(+,+,+,+)$. Such changes in signature indicate a shift from $(3+1)$ spacetime to a $4$-dimensional spatial geometry, or equivalently, a shift from the Lorentzian to the Euclidean geometry. Therefore, the CTC does not appear in the model in this range. As an extension of this work, the effects of such a non-constant scalar field can also be considered within the framework of the Brans-Dicke theories [44].
@@ 5 Conclusions
The KK theory unifies gravitation and electromagnetism by adding an extra dimension, while assuming that this extra dimension is curled to a very small scale, such as the Planck scale. After compactification in four dimensions, the KK metric leads to Einstein equations and Maxwell equations of electromagnetism. On the other hand, the Lorentzian Taub-NUT metric, as a peculiar solution to the Einstein field equations of GR in $(1+3)$ dimensions, introduces a parameter called the NUT charge that leads to a gravitomagnetic field, which originates from the gravitational magnetic monopole and does not arise in the weak-field slow-motion limit of GR. Such a monopole exhibits a feature called the Misner string, a singularity analogous to the Dirac string in electromagnetism. One way to avoid the Misner string is to make time periodic, which automatically leads to CTCs. However, the existence of CTCs challenges the causality principle and has led to the idea that Taub-NUT is more of a gravitational toy model than a realistic physical solution.
@@
In this process, we have attained gravitation and the gravitomagnetism, associated with the NUT charge, as two distinct fields. In fact, in this way, a new gravitomagnetic field has emerged from the temporal dimensional reduction. Moreover, such a compactification leads to a specific form of the dilaton field, which can have interesting effects in the framework of the Brans-Dicke theories. Remarkably, the CTC issue is resolved because for $r_{-}<r<r_{+}$ including $r=l$, the time direction becomes spacelike, changing the $(1+3)$ spacetime to a $4$-dimensional spatial geometry, and hence, the CTC does not appears in the model in this range.
We should point out that although the results have been derived from classical field equations in the Lorentzian signature, they may admit a Euclidean counterpart, subject to appropriate analytical continuation and boundary conditions, in which the NUT charge can be interpreted as a quantized topological charge, arising from the non-trivial fiber bundle structure and the regularity conditions in the Euclidean gravity [45].
@@
Action (5) as it stands, when $\phi$ is enclosed inside the parentheses, is actually a Barns-Dicke action. Hence, when varying it with respect to the $4$-dimensional metric $g_{\mu\nu}$, the scalar field $\phi$ and the vector potential $A_{\mu}$, the effect of its last (kinetic) term (i.e., $\partial^{\alpha}\phi\,\partial_{\alpha}\phi/\phi$) appears as the additional term $(g_{\alpha\beta}\partial^{\mu}\phi\,\partial_{\mu}\phi/2-\partial_{\alpha}\phi\,\partial_{\beta}\phi)/\phi^{2}$ in the first equation of Eqs. (7). Whereas, such a term does not exist there because action (5) is obtained by the dimensional reduction procedure of action (4). Indeed, for the last term inside the parentheses of action (5), one can use the relation $\partial^{\alpha}\phi\,\partial_{\alpha}\phi/\phi^{2}=\Box\phi/\phi-\nabla_{\mu}(\partial^{\mu}\phi/\phi)$. Since the second term in this relation is a complete derivative, when considered in action (4), it does not contribute in the corresponding field equations and can be ignored. Then, the remaining first term in this relation, when considered in action (5) and $\phi$ is placed inside the parentheses, is also a complete derivative and hence, does not contribute in the corresponding field equations either. Similar argument holds for action (14) and its corresponding field Eqs. (16).
Meanwhile, utilizing relations (10) and (12) and the definition of $\tilde{F}_{ab}$, and calculating the Ricci and Einstein tensors $\tilde{R}_{ab}$ and $\tilde{G}_{ab}$, and the Ricci scalar R, while using the Christoffel symbols provided in the following, it is straightforward to show that Eqs. (16) hold. For example, to prove the last equation of Eqs. (16), utilizing the standard relation

### v5

@@ Abstract
We reduce the Taub-NUT metric dimensionally to three spatial dimensions by treating time as an extra curled dimension, and end up with the $3$-dimensional Einstein field equations plus a corresponding Maxwell type equations for a gravitomagnetic field, associated with the NUT charge, which also acts as a source for the Einstein field equations. In this approach, the Taub-NUT metric can be envisaged as a $(1+3)$-dimensional analogue of the $(1+4)$-dimensional metric of the Kaluza-Klein theory. Hence, in four dimensions, it unifies gravitation and gravitomagnetism, associated with the NUT charge, in the same footing that the Kaluza-Klein theory unifies gravitation and electromagnetism in five dimensions. In fact, in this way, gravity and gravitomagnetism, associated with the NUT charge, appear as two distinct fields that emerge from the temporal dimensional reduction. We also introduce a relation between the $4$-dimensional gravitational constant and the NUT charge.
@@ 1 Introduction
On the other hand, the Taub-NUT (Newman-Tamburino-Unti) solution [18, 19] to the Einstein equations of GR has been extensively studied in the context of gravitational physics and mathematical relativity [20][22]. One the interesting aspects of this solution is the existence of the NUT charge, which is the source of the gravitomagnetic field [^4] Such a gravitomagnetic field is different from the one resulting from the weak-field and slow-motion limit of GR [23], and was experimentally confirmed in $2011$ [24]. due to the gravitomagnetic monopoles [25][27], the gravitational analogue of the magnetic monopoles [28]. However, a problem with the Taub-NUT solution is the string singularity called the Misner string. One way to eliminate the Misner string is to make the time coordinate periodic [29], which leads to the existence of closed time-like curves (CTCs). CTCs are a hallmark of the Taub-NUT solution and have profound implications for causality in GR. A CTC is a path through spacetime that loops back on itself while remaining time-like throughout.
However, CTCs are not unique to the Taub-NUT metric, but appear in several solutions of the Einstein field equations, particularly in spacetimes with unusual topology or specific symmetry properties. Notable examples include: The Gödel solution as a rotational cosmological model that admits CTCs [30] and violates global causality everywhere. Also, the interior region (inside the event horizon) of a rotating black hole, as indicated by the Kerr and Kerr-Newman metrics [31], can host CTCs, particularly near the ring singularity. These CTCs are confined to the interior of black holes and do not affect causality in the external spacetime [31]. Traversable wormholes can create CTCs if their ends experience different time dilation effects [32]. Certain configurations of rotating cosmic strings can admit CTCs due to frame-dragging effects [33]. Cylindrical distributions of rotating dust, called the van Stockum dust, can produce CTCs, depending on the density and angular velocity of the dust [34]. It is worthwhile to mention that adjusting the boundary conditions at infinity is a way to avoid CTCs [29]. Nonetheless, such an approach to eliminate CTCs in the 4-dimensional Taub-NUT metric comes with trade-offs such as restricting spatial coordinates [20].
@@ 2 Brief Review of the Kaluza-Klein Theory
$S^{\rm(KK)}\equiv\hat{S}=\frac{1}{{\kappa^{2}}}\int d^{4}x\,\sqrt{-g}\,\phi\,\left(R-\frac{1}{4}\phi^{2}\kappa^{2}F_{\alpha\beta}F^{\alpha\beta}+\frac{2}{3}\,\frac{\partial^{\alpha}\phi\,\partial_{\alpha}\phi}{\phi^{2}}\right),$ (5)
@@ 3 The Taub-NUT Solution
The main significance of the Taub-NUT solutions is the existence of the Misner string at $\theta=0$ or $\theta=\pi$, which makes the $1$-form $d\varphi$ ill-defined at either of these points. To get rid of the singularity at $\theta=0$, one can shift the time coordinate such that $t^{\prime}=t+2n\varphi$, and for the point $\theta=\pi$, one can take $t^{\prime\prime}=t-2n\varphi$, where $n$ is an integer. Since $\varphi$ is periodic by $2\pi$, one has to identify $t^{\prime}$ and $t^{\prime\prime}$ by modulo $8\pi n$, which boils down to imposing the periodicity to the time coordinate as $t\rightarrow t+8\pi n$. Doing this makes the $1$-form $d\varphi$ well-defined but leads to CTCs in the Lorentzian metric.
@@ 4 Time as an Extra Dimension in Taub-NUT Metric
In this section, we try to establish a connection between the Taub-NUT solution and the KK framework. The main purpose is to start from the Taub-NUT metric in a $4$-dimensional spacetime coordinate, when $l\neq 0$, and eventually arrive at a $3$-dimensional spatial metric plus a compactified or curled time dimension. Actually, we assume that there is a dynamical mechanism that drags the size of time dimension down and keeps it in a stable way at an unobservably tiny scale. To begin, we rewrite the $4$-dimensional Taub-NUT metric (9) in the matrix form [^9] We represent $3$-dimensional quantities with a tilde, $\widetilde{}$ , over them.
@@
where the lowercase Latin indices run from one to three, the solution to the $3$-dimensional vector potential is obtained as $\tilde{\bf A}=(0,0,\cos{\theta})$, the metric signature is (,+,+,+)(-,+,+,+) provided that $\tilde{g}_{33}-4l^{2}\Phi^{2}(r)\cos^{2}{\theta}>0$, and $g=-\Phi^{2}(r)\,\tilde{g}$, where $\tilde{g}$ is the determinant of the $3$-dimensional metric. Moreover, in the range $r<r_{-}$ and $r>r_{+}$, the solution to the $3$-dimensional metric is
@@
with the signature $(+,+,+)$ in the spherical coordinates. However, in the range $r_{-}<r<r_{+}$ including $r=l$, the scalar field $\Phi(r)$ (10) becomes imaginary and hence the signature of metric (12) changes to (,+,+)(-,+,+). Indeed, in these ranges, in metric (9), time represents a spacelike coordinate and $r$ a timelike one.
@@
where $\oint dx^{0}$} is the result of integration along the $x^{0}$-direction. The reduced action involves the gauge field $\tilde{A}_{a}$, which is a gravitomagnetic type gauge field associated with the NUT charge, and $\Phi(r)$, which plays the role of a scalar or a dilaton field within the context of the KK theory.
@@ 5 Conclusions
The KK theory unifies gravitation and electromagnetism by adding an extra dimension, while assuming that this extra dimension is curled to a very small scale, such as the Planck scale. After compactification in four dimensions, the KK metric leads to Einstein equations and Maxwell equations of electromagnetism. On the other hand, the Lorentzian Taub-NUT metric, as a peculiar solution to the Einstein field equations of GR in $(1+3)$ dimensions, introduces a parameter called the NUT charge that leads to a gravitomagnetic field, which originates from the gravitational magnetic monopole and does not arise in the weak-field slow-motion limit of GR. Such a monopole exhibits a feature called the Misner string, a singularity analogous to the Dirac string in electromagnetism. One way to avoid the Misner string is to make time periodic, which automatically leads to CTCs.
@@
In this process, we have attained gravitation and the gravitomagnetism, associated with the NUT charge, as two distinct fields. In fact, in this way, a new gravitomagnetic field has emerged from the temporal dimensional reduction. Also, we have analogously obtained a relation between the $4$-dimensional Newtonian gravitational constant and the NUT charge.
@@
Utilizing relations (10) and (12) and the definition of $\tilde{F}_{ab}$, and calculating the Ricci and Einstein tensors $\tilde{R}_{ab}$ and $\tilde{G}_{ab}$, and the Ricci scalar R, while using the Christoffel symbols provided in the following, it is straightforward to show that Eqs. (16) hold. For example, to prove the last equation of Eqs. (16), utilizing the standard relation
@@
It is worthwhile to mention that action (5) as it stands, when $\phi$ is enclosed inside the parentheses, is actually a Brans-Dicke action. Hence, when varying it with respect to the $4$-dimensional metric $g_{\mu\nu}$, the scalar field $\phi$ and the vector potential $A_{\mu}$, the effect of its last (kinetic) term (i.e., $\partial^{\alpha}\phi\,\partial_{\alpha}\phi/\phi$) appears as the additional term $(g_{\alpha\beta}\partial^{\mu}\phi\,\partial_{\mu}\phi/2-\partial_{\alpha}\phi\,\partial_{\beta}\phi)/\phi^{2}$ in the first equation of Eqs. (7). Whereas, such a term does not exist there because action (5) is obtained by the dimensional reduction procedure of action (4). Indeed, for the last term inside the parentheses of action (5), one can use the relation $\partial^{\alpha}\phi\,\partial_{\alpha}\phi/\phi^{2}=\Box\phi/\phi-\nabla_{\mu}(\partial^{\mu}\phi/\phi)$. Since the second term in the right hand-side of this relation is a complete derivative, when considered in action (4), it does not contribute in the corresponding field equations and can be ignored. Then, the remaining first term in the right hand-side of this relation, when considered in action (5) and $\phi$ is placed inside the parentheses, is also a complete derivative and hence, does not contribute in the corresponding field equations either. Similar argument holds for action (14) and its corresponding field Eqs. (16).